### v1

# Diagnosing Quantum Many-body Chaos in Non-Hermitian Quantum Spin Chain via Krylov Complexity

Abstract
We investigate the phase transitions from chaotic to non-chaotic dynamics in a quantum spin chain with a local non-Hermitian disorder, which can be realized with a Rydberg atom array setting. As the disorder strength increases, the emergence of non-chaotic dynamics is qualitatively captured through the suppressed growth of Krylov complexity, and quantitatively identified through the reciprocity breaking of Krylov space. We further find that the localization in Krylov space generates another transition in the weak disorder regime, suggesting a weak ergodicity breaking. Our results closely align with conventional methods, such as the entanglement entropy and complex level spacing statistics, and pave the way to explore non-Hermitian phase transitions using Krylov complexity and associated metrics.
Introduction.— Diagnosis of chaos emerging in quantum systems attracts a growing interest, with a variety of measures including Lyapunov exponents [1, 2, 3], out-of-time-order correlations [4, 5, 6], Loschmidt echo [7, 8, 9], as well as spectral characteristics revealed through level spacing distributions [10, 11, 12, 13, 14] and spectral form factors [15, 16, 17]. Recently, Krylov complexity has emerged as an innovative approach to the study of quantum chaos [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Krylov complexity quantifies the spreading of quantum operators [18] or states  [20] in Krylov space, capturing key features of the underlying dynamics.
Figure 1: (a) Disordered non-Hermitian XY model, with XY coupling $J$, transverse field $h$, site-dependent detuning $\Delta_{j}\in[-W_{\Delta},W_{\Delta}]$, and dissipation $\gamma_{j}\in[-W_{\gamma},W_{\gamma}]$. (b) Tunneling rates $b_{j}$ and $c_{j}$ are not conjugate for non-Hermitian systems. (c) Sketch of the phase diagram with increasing $W_{\gamma}$. (d) Krylov wave functions in chaotic regimes with delocalized (left) and localized (middle) profiles, and non-chaotic (right) regimes, at mid-term $t=300$ (solid blue) and long-term $t=1000$ (dashed red). See text for details.
A pivotal insight in this field is the operator growth hypothesis [18], which suggests that information scrambling and chaos can be identified by observing exponential growth in Krylov complexity at early times. The exponential growth is attributed to the linearity of Lanczos coefficients, which govern the tunneling rates in Krylov space. In contrast, another hypothesis focuses on the long-term dynamics and proposes a Tsallis $q$-log behavior of the Lanczos coefficients throughout the entire Krylov space [24, 32], where chaos can be signified by the linear growth in complexity at intermediate times, followed by a peak before saturation [20, 24, 27]. Additionally, the fluctuation in Lanczos coefficients, known as Krylov variance, has been shown to correlate with the delocalization of the Krylov wave function and, consequently, with ergodic behaviors [19, 21, 25, 29, 28, 33]. Despite extensive investigations in the Hermitian regime, challenges emerge when extending to open quantum systems. While Krylov complexity appears to be significantly reduced under dissipation [34, 35, 36, 37, 38, 39], the Lanczos coefficients deviate from the established hypotheses for Hermitian systems, particularly when the absence of Hermiticity of Krylov space is overlooked. Furthermore, the critical behavior and universal properties near phase transitions in open systems remain largely unexplored.
In this work, we investigate Krylov complexity of a non-Hermitian XY model with random local gain and loss. We show that the evolution of Krylov wave functions and complexity growth are indicative of phase transitions of quantum chaos. Through a closer examination of Lanczos coefficients, we uncover that the Krylov variance elucidates critical points and universal features for localization of Krylov wave function which indicates a weak ergodicity breaking. We introduce a metric of reciprocity to underscore the suppression of quantum chaos, where the onset of non-reciprocity marks the corresponding critical point. The critical points determined using our method are quantitatively consistent with established metrics, including entanglement entropy variance [40, 41] and complex level spacing statistics [12, 13, 14]. Our framework is also applicable to other ergodicity-breaking phenomena, including many-body localization [21, 42], quantum scars [43], and Hilbert space fragmentation [44], and holds potential for extensions into machine learning applications [45]. The non-Hermitian XY model can be realized with Rydberg atom array quantum simulators. Our work thus opens avenues to experimentally explore complexities with observables such as Loschmidt echo [7, 46].
Model.— We study a one-dimensional disordered XY spin chain under both transverse and longitudinal fields, as depicted in Fig. 1(a). The Hamiltonian is given by
$H=\sum_{j=1}^{L-1}J\left({\hat{\sigma}_{j}^{x}\hat{\sigma}_{j+1}^{x}+\hat{\sigma}_{j}^{y}\hat{\sigma}_{j+1}^{y}}\right)+\sum_{j=1}^{L}{h\hat{\sigma}_{j}^{x}+\mathcal{D}_{j}\hat{\sigma}_{j}^{z}},$ (1)
where $\hat{\sigma}_{j}^{x}$, $\hat{\sigma}_{j}^{y}$, $\hat{\sigma}_{j}^{z}$ are Pauli matrices acting on the $j$-th site. The longitudinal field is random and complex, $\mathcal{D}_{j}=\Delta_{j}+{\rm i}\gamma_{j}$, where the real part $\Delta_{j}$ and imaginary part $\gamma_{j}$ give the detuning, and the gain or loss, respectively. The parameters $\Delta_{j}$ and $\gamma_{j}$ are independently and uniformly distributed within $[-W_{\Delta},W_{\Delta}]$ and $[-W_{\gamma},W_{\gamma}]$, correspondingly. In the following calculations, we scale the Hamiltonian with the XY interaction strength $J$. To explore the non-Hermitian effects, we vary $W_{\gamma}$ while fixing $h=0.5$ and $W_{\Delta}=1$ throughout the work. The model can be realized with Rydberg atom arrays, which will be discussed later.
In the Krylov space, the dynamics are mapped to a single-particle tight-binding model [see Fig. 1(b)]. The non-Hermitian Hamiltonian displays vastly different dynamics as the disorder strength $W_{\gamma}$ increases. Specifically, two phase transitions are observed, where a suppression of spreading within finite Krylov subspace takes place at $W_{\gamma}=W_{L}\approx 0.0351$ revealing a localization feature, and a transition from chaotic to non-chaotic dynamics at $W_{\gamma}=W_{C}\approx 1.647$. The phase diagram is depicted in Fig. 1(c). These phases can be captured by the spreading of the Krylov wave function, illustrated in Fig. 1(d) [see Supplemental Materials (SM) for more examples]. For $W_{\gamma}<W_{L}$, the Krylov wave function is fully delocalized across the entire Krylov space at late times ($t=1000$). For $W_{\gamma}>W_{L}$, conversely, a weak breaking of ergodicity may be indicated by the localized Krylov wave function with exponentially small occupation for $n\gtrsim\mathcal{O}(L^{2})$. In the non-chaotic regime, $W_{\gamma}<W_{C}$, the wave function is further confined below $n\lesssim\mathcal{O}(L)$. We also observe prethermal behavior in the Krylov delocalized regime, where the wave function is confined near $n\approx 0$ at mid-term times ($t=300$), contrasting with other phases. Krylov complexity and Lanczos coefficients capture critical behaviors around the phase transition points. This allows us to determine the critical points through a number of measures, which will be discussed in detail in the following.
Figure 2: (a) Krylov complexity growth. With $W_{\Delta}=0.5$, $W_{\gamma}$ varies from $0.001$ to $5$ across system sizes (a1) $L=6$, (a2) $L=8$, and (a3) $L=10$. Dashed lines are the quadratic ($\sim t^{2}$) and dash-dot lines the linear (tsimilar-$\sim t$) trends, respectively. (b) Lanczos coefficients for $L=10$ and $W_{\gamma}=$ (b1) 0.0118, (b2) 0.2, (b3) 1.0, and (b4) 3.0. Upper panels show $\langle|b_{n}|\rangle$ (solid blue) and $\langle|c_{n}|\rangle$ (dashed red), with dotted lines representing $\sim n$ for $n<L$, and $\sim\sqrt{1-n/2^{L}}$ for $n>L$. Lower panels show $\langle\cos\theta_{n}\rangle$. (c) The Krylov variance given by Eq. (4), and the inset is the finite-size scaling, from which we obtain $W_{L}=0.0351$. (d) Krylov reciprocity given by Eq. (5). The inset is the finite-size scaling yielding $W_{C}=1.647$. Simulations are sampled and averaged over 1000, 600, and 400 trajectories for $L=6,8$, and $10$, respectively.
Krylov complexity.— The mathematical framework for representing quantum models in Krylov space involves a tridiagonalization process [47]. Krylov space is constructed by iteratively applying the Hamiltonian to an initial state vector and orthogonalizing it against the preceding basis vectors. This process is efficiently implemented using the bi-Lanczos algorithm, which is capable of dealing with non-Hermitian systems [36, 38, 37]. In Krylov space, the evolution of the quantum system is encapsulated as an effective single-particle tight-binding model on a semi-infinite chain [18]. The Krylov wave function $\varphi_{n}$ is governed by the discrete equation (see SM for details),
$i\dot{\varphi}_{n}=b_{n}\varphi_{n-1}+a_{n}\varphi_{n}+c_{n+1}\varphi_{n+1},$ (2)
where $\varphi_{n}$ represents the distribution over the Krylov space. The Lanczos coefficients $b_{n}$ and $c_{n}$ are generally not complex conjugates, due to the non-Hermiticity of the Hamiltonian. The initial state is $\varphi_{0}=1$ and $\varphi_{n}=0$ for $n>0$.
Krylov complexity quantifies the spreading of the Krylov wave function by its center of mass, defined as
$C_{K}=\sum_{n}n\left|\varphi_{n}(t)\right|^{2}.$ (3)
The growth of Krylov complexity is a rich, multistage phenomenon. Figures 2(a1-a3) illustrate the complexity growth for system sizes $L$ = 6, 8, and 10 with the initial state $\left|\psi_{0}\right\rangle=\left|+\right\rangle^{\otimes L}$, where $\left|+\right\rangle=(\left|\uparrow\right\rangle+\left|\downarrow\right\rangle)/\sqrt{2}$. In case of Hermitian systems ($W_{\gamma}=0$), the Krylov complexity typically shows generic quadratic-linear-peak-saturate shape, reminiscent of the slope-dip-ramp-plateau profile of the spectral form factor [24]. In our non-Hermitian spin chain model, we find distinctive features beyond these trends, indicating Krylov localization and chaos-non-chaos transition as $W_{\gamma}$ increases.
The short-term growth of the Krylov complexity is universally quadratic, $C_{K}(t)\sim(1+W_{\gamma}^{2})t^{2}$ (see SM). The absence of exponential growth suggests a deviation from the operator growth hypothesis, as also corroborated by the nonlinear behavior of Lanczos coefficients [Fig. 2(b1-b4)]. Thus, it precludes chaos diagnosis using the early time dynamics in our model.
On the intermediate time scale, the Krylov complexity grows nonlinearly, deviating from the linear growth associated with spectral rigidity reported in previous studies [24]. Notably, for $W_{\gamma}<W_{C}$, $C_{K}$ is bounded by a linear envelope, and its long-term saturation value increases as $W_{\gamma}$ decreases. These phenomena indicate that chaos suppression correlates with confined spreading of Krylov wave function [Fig. 1(d)]. In chaotic regimes, the Krylov wave functions extend beyond the system size, as a result of the involvement of many excited states. Ref. [24] suggests a linear ramp of the Krylov complexity due to the Lanczos coefficients following $b_{n}\propto\sqrt{1-n/2^{L}}$, which is a special case of Tsallis $q$-log statistics with $q=0$ [32]. In contrast, we find $q>0$ in our results (see SM), as shown in Figs. 2(b1-b4). The mid-term ramp, on the other hand, indicates spectral rigidity as it is constrained by a linear envelope [Fig. 2(a)].
In the Krylov delocalized regime ($W_{\gamma}<W_{L}$), the Krylov complexity shows prethermal behaviors, where oscillations and a low saturation plateau emerge before a rapid increase to higher values. This behavior stems from dynamics confined to a small Krylov subspace due to small overlaps between the initial state and eigenstates with large imaginary components. For example, for $W_{\gamma}=0.022$ and at $t=300$, the system forms a prethermal state where the Krylov wave function is localized at $n<\mathcal{O}(L)$ [Fig. 1(d)]. At longer times $t\gtrsim 10^{3}$, $C_{K}$ saturates as the Krylov wave function completely delocalizes. As system size increases, $C_{K}$ exhibits damped oscillations, and the saturation time grows significantly.
Krylov variance and reciprocity.— Beyond the growth of $C_{K}$, a metric known as Krylov variance [22] has been proposed to measure the fluctuation of the Lanczos coefficients, which characterizes the localization of the Krylov wave function as an analogue to Anderson localization [48]. In Hermitian systems, the Krylov variance is defined as $\sigma_{K}^{2}={\rm Var}\{\ln|b_{2n-1}/b_{2n}|\}$ [22]. Similarly, we define the variance of the non-Hermitian system by replacing $b_{n}$ with $j_{n}=|b_{n}c_{n}|^{1/2}$, yielding
$\sigma_{K}^{2}={\rm Var}\left\{\ln\frac{j_{2n-1}}{j_{2n}}\right\}\quad,\quad n=1,2,3,\cdots.$ (4)
Figure 2(c) shows the Krylov variance first grows with increasing $W_{\gamma}$ and saturates at larger $W_{\gamma}$. Through finite-size scaling, we find $\sigma_{K}^{2}$ can be universally fitted with function $y=\sigma_{K}^{2}L^{-\beta}$ against $x=(W_{\gamma}-W_{L})L^{\alpha}$ [see inset of Fig. 2(c)]. This allows us to identify a critical point, $W_{L}=0.0351$, with exponents $\alpha=1.816$ and $\beta=0.112$. When $W_{\gamma}<W_{L}$, the Krylov variance is small, which results in the delocalization of the Krylov wave function [Fig. 1(d)]. When $W_{\gamma}>W_{L}$, the Krylov wave function localizes in a finite region. Here, the system is still chaotic, as supported by the level repulsion (as discussed below). Yet, the localized Krylov wave function and hence the suppression of Krylov complexity emerge as a signature of ergodicity breaking [21, 29, 33].
In non-Hermitian systems, Lanczos coefficients $b_{n}$ and $c_{n}$ are not complex conjugates, violating the reciprocity of the Krylov space. To quantify this, we evaluate arguments of the Lanczos coefficients, $\theta_{n}=\arg(b_{n}c_{n})$, such that $\cos\theta_{n}=1$ represents reciprocal tunneling, and $\cos\theta_{n}=-1$ gives the maximally non-reciprocal tunneling, which is associated with rapid relaxation. Figures 2(b1-b4) show the profiles of $\cos\theta_{n}$ as $W_{\gamma}$ increases. Specifically, for $W_{\gamma}<W_{L}$, the entire Krylov space remains reciprocal, supporting delocalization of the Krylov wave function. For $W_{L}<W_{\gamma}<W_{C}$, tunneling within $n<L$ is reciprocal, but a rapid relaxation occurs as the Krylov wave function spreads to $n>L$. For $W_{\gamma}>W_{C}$, the Krylov wave function rapidly relaxes within $n<L$, as the first few sites become maximally non-reciprocal.
To leverage the sign-flip behavior of $\cos\theta_{n}$, we introduce a parameterized metric,
$R_{K}^{(d)}=\frac{1}{d}\sum_{n=1}^{d}\cos\theta_{n},$ (5)
which is insensitive to the choice of $d$ for a certain range smaller than $L$. We show this reciprocal metric in Fig. 2(d) for $d=4$, and more examples in SM. For different $L$, $R_{K}^{(d)}$ collapses to $y=R_{K}^{(d)}$ with $x=(W_{\gamma}-W_{C})L^{\alpha}$, yielding the mean critical point $W_{C}=1.647$ and exponent $\alpha=0.820$ for $d\in\{4,5,6\}$, with the relative deviation $<1\%$. The universal scaling thus shows the transition from the chaotic to non-chaotic phases.
[…]
Entanglement entropy and level spacing.— To provide a comprehensive view of the aforementioned phase transitions, we bridge the findings from the Krylov complexity with conventional signatures of quantum chaos including entanglement entropy and complex level spacing. It is known that the chaos-non-chaos transition can be signified by the divergence of the variance of the bipartite entanglement entropy [40]. We show the standard deviation of the bipartite entanglement entropy, $\sigma_{S}$, in Fig. 3(a). A divergent peak is observed as $L\to\infty$. The data can be universally described by function $y=\sigma_{S}L^{-\beta}$ with respect to $x=(W_{\gamma}-W^{\prime}_{C})L^{\alpha}$ [inset of Fig. 3(a)]. This leads to the critical point $W^{\prime}_{C}=1.763$ with exponents $\alpha=2.013$ and $\beta=1.228$. Furthermore, a critical behavior at very small $W_{\gamma}$ can be seen in Fig. 3(b). Using a similar scaling method, we identify the critical point $W^{\prime}_{L}=0.0409$ and exponents $\alpha=1.514$ and $\beta=-0.054$. The critical points extracted from $\sigma_{S}$ are similar to the ones of the Krylov complexity. In addition, we obtain a different critical point at $W^{*}=3.236$ with the finite-size analysis of the entanglement entropy, $S$ (see SM). This attributes to the transition to integrability as the system is well described by product states when $W_{\gamma}$ is sufficiently large.
[…]
Conclusions.— In conclusion, we have studied non-Hermitian quantum chaos in the disordered XY model, revealing two critical phase transitions including the chaotic-non-chaotic transition and Krylov localization. Through examing the Krylov complexity and Lanczos coefficients characterized by the Krylov variance and reciprocity, we identify the critical points and exponents of these phase transitions. Our results are corroborated by established measures including entanglement entropy and complex level spacings, which showcase the applicability of Krylov complexity and the associated measures in the study of quantum chaos. The sign-flip of Krylov reciprocity provides a quantitatively more precise criterion for chaos suppression than the entanglement entropy. Furthermore, we observe intricate chaotic behavior in non-Hermitian systems, where prethermalization may exist in the Krylov delocalized regime.
The non-Hermitian spin model can be realized in Rydberg atom arrays, where spin states are encoded by two Rydberg states [50]. The XY interactions correspond to the dipole-dipole interactions between the Rydberg states, and the dissipation can be induced by coupling the Rydberg states to low-lying electronic states [51]. Randomness of the local fields can be introduced by AC-Stark shifts with addressing lasers. In generic quantum simulators, non-Hermiticity can be realized through Sz.-Nagy dilation [52, 53] and linear combination of unitaries [54, 55]. Finally, to measure the Krylov complexity, it has been proposed to apply measurable bases for estimation [56], and the measure of Loschmidt echo is potentially related to the ergodicity breaking [Fig. 1(d)].
Acknowledgements.
We are grateful to Dong-Ling Deng, Dong Yuan, Si Jiang, Chang Liu (SQZI), Chang Liu (NUS), Tianyi Yan, Xiaopeng Li, Xingze Qiu and Saud Čindrak for fruitful discussions. Y.Z. acknowledges support from the Shanghai Qi Zhi Institute Innovation Program SQZ202317. W.X. acknowledges support from the National Natural Science Foundation of China/Hong Kong RGC Collaborative Research Scheme (Project CRS CUHK401/22) and New Cornerstone Science Foundation. L.L. acknowledges support from the National Key Research and Development Program of China (Grant No. 2021YFA1402003) and the National Natural Science Foundation of China (Grant Nos. 12374329 and U21A6006). W.L. acknowledges support from the EPSRC through Grant No. EP/W015641/1, the Going Global Partnerships Programme of the British Council (Contract No. IND/CONT/G/22-23/26), and the International Research Collaboration Fund of the University of Nottingham.